A point has its centre (2,-2). The radical axis of this circle with the circle 3(x^2 +y^2) - 6x +9y +1=0

Grade:11

1 Answers

Harsh Patodia IIT Roorkee
6 years ago
radical axis is s1-s2=0 here S1=(x-2)2+(y+2)2=x2+y2-4x+4y+8 and S2=x2+y2-2x+3y+1/3=0
S1-S2=-2x+y+8-1/3=-6x+3y+23
=6x-3y-23=0